# Non-linear degenerate parabolic flow equations and a finer differential structure on Wasserstein spaces

### v1

@@ 3.7.2 Grönwall type estimates and Hausdorff–continuity of the set of solutions
Let us first work with $(M,g)=(\mathbb{R}^{n},g^{eucl})$. Let $t\in[0,T]$ and $\varepsilon>0$. Denote for all $N\in\mathbb{N}$, $(x_{N,i})\coloneqq\mathcal{}\mathfrak{P}(\frac{1}{N})$ in increasing order. Note that, for all $t\in[0,T]$, $\mathcal{W}_{2}(V(t),W(t))\leq 2R$. Hence, (3.60) in Lemma 3.31 yields that there exists $N\in\mathbb{N}$—depending on $\mathcal{W}_{2}^{\infty}(V,W)$—, such that for all $n\geq\mathbb{N}$,
[…]
Now, let $i\in I_{n}$ and suppose that there exists $C>0$ such that,
$\mathcal{W}_{2}\bigl{(}\mu^{\mathcal{}\mathfrak{P}\bigl{(}\frac{1}{N}\bigr{)}}(x_{n,i}),\nu^{\mathcal{}\mathfrak{P}\bigl{(}\frac{1}{N}\bigr{)}}(x_{n,i})\bigr{)}^{2}\leq C^{2}\mathcal{W}_{2}^{\infty}(V,W)^{2}.$
Then,
$\mathcal{W}_{2}(f^{V(x_{n,i})}_{x_{n,i+1}-x_{n,i}}(\mu^{\mathcal{}\mathfrak{P}\bigl{(}\frac{1}{N}\bigr{)}}(x_{n,i})),f^{V(x_{n,i})}_{x_{n,i+1}-x_{n,i}}(\nu^{\mathcal{}\mathfrak{P}\bigl{(}\frac{1}{N}\bigr{)}}(x_{n,i})))^{2}$ (3.139)
$\qquad\leq\biggl{(}\mathcal{W}_{2}\bigl{(}\mu^{\mathcal{}\mathfrak{P}\bigl{(}\frac{1}{N}\bigr{)}}(x_{n,i}),\nu^{\mathcal{}\mathfrak{P}\bigl{(}\frac{1}{N}\bigr{)}}(x_{n,i})\bigr{)}^{2}$
$\qquad+2(x_{n,i+1}-x_{n,i})\biggl{(}9L^{2}\mathcal{W}_{2}^{\infty}(V,W)^{2}+3LRC\mathcal{W}_{2}^{\infty}(V,W)^{2}\biggr{)}$
$\qquad+4R^{2}\biggl{(}RC\mathcal{W}_{2}^{\infty}(V,W)+3L\mathcal{W}_{2}^{\infty}(V,W)\biggr{)}\frac{(x_{n,i+1}-x_{n,i})^{\frac{3}{2}}}{6}$
$\qquad+\frac{2R^{4}}{3}(x_{n,i+1}-x_{n,i})^{2}\biggr{)}e^{8R^{2}(x_{n,i+1}-x_{n,i})}.$
In particular, for all $C^{2}\leq\mathcal{B}$ for some $\mathcal{B}>0$, there exists $M\geq N$—depending on $\mathcal{W}_{2}^{\infty}(V,W)$—such that for all $n\geq M$,
$\mathcal{W}_{2}(f^{V(x_{n,i})}_{x_{n,i+1}-x_{n,i}}(\mu^{\mathcal{}\mathfrak{P}\bigl{(}\frac{1}{N}\bigr{)}}(x_{n,i})),f^{V(x_{n,i})}_{x_{n,i+1}-x_{n,i}}(\nu^{\mathcal{}\mathfrak{P}\bigl{(}\frac{1}{N}\bigr{)}}(x_{n,i})))^{2}$ (3.140)
$\qquad\leq\biggl{(}C^{2}+2(x_{n,i+1}-x_{n,i})\bigl{(}9L^{2}+3LRC+8R^{2}C^{2}+\varepsilon\bigr{)}\biggr{)}\mathcal{W}_{2}^{\infty}(V,W)^{2}$
$\qquad\leq\biggl{(}C^{2}+2(x_{n,i+1}-x_{n,i})\bigl{(}9L^{2}+3LR\sqrt{\mathcal{B}}+8R^{2}\mathcal{B}+\varepsilon\bigr{)}\biggr{)}\mathcal{W}_{2}^{\infty}(V,W)^{2}.$
Now, (3.136) follows by induction. The result for an arbitrary manifold with bounded geometry follows by a straightforward application of Nash’s embedding theorem.
For the second part, denote by $t\mapsto\mu(t)$ a smooth solution of (3.125) for $V$ in $\mathcal{W}_{p}(M)$. Hence, there exists a sequence of partitions $\biggl{(}\mathcal{}\mathfrak{P}(\frac{1}{N})\biggr{)}_{N\in\mathbb{N}}$ such that $t\mapsto\mu(t)$ is a uniform limit of $\biggl{(}\mu^{\mathcal{}\mathfrak{P}(\frac{1}{N})}\biggr{)}_{N\in\mathbb{N}}$. Now, Theorem 3.42 yields that there exists an increasing sequence $N_{k}$ such that $\biggl{(}\nu^{\mathcal{}\mathfrak{P}(\frac{1}{N_{k}})}\biggr{)}_{N\in\mathbb{N}}$ converges uniformly to a solution $t\mapsto\nu(t)$ of (3.125) for $W$ in $\mathcal{W}_{p}(M)$. The uniform convergence implies for all $t\in[0,T]$,
$\mathcal{W}_{2}(\mu(t),\nu(t))\leq\mathcal{K}\sqrt{t}\mathcal{W}_{2}^{\infty}(V,W).$ (3.141)
In particular,
$\mathcal{W}_{2}(\mu(t),\nu(t))\leq\mathcal{K}\sqrt{T}\mathcal{W}_{2}^{\infty}(V,W).$ (3.142)
Thus, for every smooth solution of (3.125) for $V$, there exists a smooth solution of (3.125) for $W$, such that
$\mathcal{W}_{2}^{\infty}(\mu,\nu)\leq\mathcal{K}\sqrt{T}\mathcal{W}_{2}^{\infty}(V,W),$ (3.143)
and vice–versa. This yields the second result. $\square$
3.7.3 Strong Generalized Central Limit Theorem and uniqueness of smooth solutions
Now, we are ready to prove our third and last main result, the uniqueness of smooth solutions of (3.14). First, we prove that the smooth solutions (3.125) are unique for sufficiently regular initial values. Then, we combine Theorem 3.43 and some of our Grönwall type estimates to conclude.
Proposition 3.48.
Let $R\geq 1$, $p\geq 2$, $\mu_{0}\in\mathcal{C}_{c}^{\infty}(M)$, and $V:[0,T]\mapsto\mathcal{W}_{2}(\mathcal{X}^{2,\infty}(M))$ be $\frac{1}{p}$–Hölder–continuous and uniformly supported in the centered ball with radius $R$. Then there exists a unique smooth solution of equation (3.125) for $V$ with initial value $\mu_{0}$ in $\mathcal{W}_{2}(M)$.
Proof.
Let $\varepsilon>0$. First, Theorem 3.25 yields that there exist $W:[0,T]\mapsto\mathcal{W}_{2}(\mathcal{X}^{2,\infty}(M))$, which is $\frac{1}{p}$–Hölder–continuous and uniformly supported in the centered ball with radius $R$ and $\lambda>0$, such that for all $t\in[0,T]$, $\square^{W(t)}$ is $\lambda$–elliptic and $\mathcal{W}_{p}^{\infty}(V,W)\leq\varepsilon$. Since the $W(t)$ are uniformly supported in the ball with radius $R$, one readily shows that for all $i,j\in\llbracket 0,n-1\rrbracket$,
@@ 3.7.3 Strong Generalized Central Limit Theorem and uniqueness of smooth solutions
are uniformly in $W^{2,\infty}(M)$ for fixed $t$ and uniformly in $\mathcal{C}^{0,\frac{1}{p}}([0,T],\mathbb{R})$ for fixed $x$. Thus, (3.125) has a unique solution $s$ for $W$ with initial value $\mu_{0}$ in $\mathcal{W}_{2}(M)$. Now $L:=\tfrac{\partial}{\partial t}-\square^{W(t)}$ is a second-order parabolic operator on $M\times(0,T]$, which is uniformly parabolic and the coefficients are uniformly $W^{2,\infty}$ in space and uniformly $\tfrac{1}{p}$–Hölder in time, so they are in $C^{2\alpha,\alpha}_{loc}(M\times(0,T])$ for $\alpha=\tfrac{1}{p}$. If $u\in\dot{\mathcal{B}}^{\prime}(M\times(0,T])$ ([16] where the decay is only spatially but uniformly in $[t,T]$ for all $t>0$) satisfies $Lu=0$ in the sense of distributions and $u(0)=\mu_{0}$. Then, by the parabolic Schauder theory, one has
@@
Theorem 3.49 (Generalized Central Limit Theorem—Strong version).
Let $p>2$ and $T>0$. Let $\mu_{0},\nu_{0}\in\mathcal{W}_{p}(M)$. Let $\bigl{(}\mathcal{}\mathfrak{P}(\frac{1}{N})\bigr{)}_{N\in\mathbb{N}}$ be a family of partitions of $[0,T]$. Assume:
@@
  • $V$ is a vector–field probability which satisfies:
    • There exists $R>0$ such that for all $\mu\in\mathcal{W}_{p}(M)$, $V[\mu]$ is supported in the centered ball with radius $R$.
@@
Then, (3.14) has a unique smooth solution $\mu:[0,T]\to\mathcal{W}_{p}(M)$ (see Definiton 3.28) and the Average Flow Approximation Series $\bigl{(}\mu^{\mathcal{}\mathfrak{P}(\frac{1}{N})}\bigr{)}_{N\in\mathbb{N}}$ (see Definition 3.27) converges uniformly to $t\mapsto\mu(t)$ in $\mathcal{W}_{p}(M)$. Furthermore, $t\mapsto\mu(t)$ is $\frac{1}{p}$–Hölder–continuous.
@@
Let $\mu_{0}\in\mathcal{C}^{\infty}_{c}(M)$. Proposition 3.48 yields that for all $\frac{1}{p}$–Hölder–continuous $\mu:[0,T]\to\mathcal{W}_{p}(M)$ with $\mu(0)=\mu_{0}$, the linear MDE (3.125) has a unique smooth solution for $t\mapsto V[\mu(t)]$ with initial value $\mu_{0}$ in $\mathcal{W}_{p}(M)$. Hence, Theorem 3.43 yields that (3.14) has a unique smooth solution with initial value $\mu_{0}$ in $\mathcal{W}_{p}(M)$. Finally, we extend the uniqueness on arbitrary $\mu_{0}\in\mathcal{W}_{p}(M)$ with the last point of Theorem 3.44. Now, the rest of the statement is a consequence of Theorem 3.35. $\square$
@@ 3.7.4 A counter–example to general uniqueness of solutions
General solutions of 3.14 are however not unique in general, even if $V$ is Lipschitz and the $\square^{V}_{\mu}$ are uniformly elliptic for all $\mu\in\mathcal{W}_{p}(M)$. Let us give an example of two distinct solutions to (3.14) with same initial value.
@@
There exists a vector–field probability $V$ which is Lipschitz with respect to the pp–Wasserstein distance and such that $\square^{V}_{\mu}$ is strongly elliptic for all $\mu\in\mathcal{W}_{p}(M)$, such that solutions to 3.14 are not unique.
@@
For any $r\in\mathbb{R}$ and $v_{1},v_{2}\in\mathbb{R}^{2}$ with $\bigl{<}v_{1},v_{2}\bigr{>}=0$, denote by $\mathcal{C}_{r}$ the cercle with radius $r$ and by $\mathcal{E}_{v_{1},v_{2}}$ the ellipse with semimajor axis $v_{1}$ and semiminor axis $v_{2}$. Finally, define the vector–field–probability
$V:\mathcal{W}_{p}(\mathbb{R}^{2})$ $\mapsto\mathcal{W}_{p}(\mathcal{X}^{2,\infty}(\mathbb{R}^{2}))$ (3.149)
@@
where, by abuse of notation, for all $X\in\mathbb{R}^{2}$, we identify the the constant vector–field $v_{X}:y\mapsto X$ with $X$ here. Now, we show that $V$ is locally Lipschitz with respect to the pp–Wasserstein distance around $\mu_{0}$, where $\mu_{0}$ is the standard normal distribution. Using Proposition 2.8, there exist $R>0$, $B>0$ and $b>0$, such that for all $\mu\in B(\mu_{0},R)$, $r({\operatorname{Cov}}(\mu))\geq b$ and $\rho({\operatorname{Cov}}(\mu))\leq B$. Now, there exist $C>0$, such that for all $\rho,r_{1},r_{2}\in[b,B]$, $\bigl{|}\Delta_{\rho}(r_{1})-\Delta_{\rho}(r_{2})\bigr{|}\leq C|r_{1}-r_{2}|$. Now, Proposition 2.8 yields that $V$ is locally Lipschitz around $\mu_{0}$. Furthermore, $V$ is uniformly bounded by $M>0$.
Now, we show that there exist to distinct solutions to (3.14) with initial value $\mu_{0}$. Let $T>0$. Define
$\mu_{1}:[0,T]\mapsto\biggl{(}(x,y)\mapsto\frac{1}{2\pi\sqrt{(t+1)e^{t}}}e^{-\frac{1}{2}\bigl{(}\frac{x^{2}}{t+1}+\frac{y^{2}}{e^{t}}\bigr{)}}\biggr{)}$ (3.150)
$\mu_{2}:[0,T]\mapsto\biggl{(}(x,y)\mapsto\frac{1}{2\pi\sqrt{(t+1)e^{t}}}e^{-\frac{1}{2}\bigl{(}\frac{y^{2}}{t+1}+\frac{x^{2}}{e^{t}}\bigr{)}}\biggr{)}.$
@@
$\frac{\operatorname{d}}{\operatorname{d}t}\mu_{1}(t)={\operatorname{Cov}}(\mu_{1}(t))_{1,1}\frac{\operatorname{d}^{2}}{\operatorname{d}x^{2}}\mu_{1}(t)+\frac{\operatorname{d}^{2}}{\operatorname{d}y^{2}}\mu_{1}(t),$ (3.151)
$\frac{\operatorname{d}}{\operatorname{d}t}\mu_{2}(t)={\operatorname{Cov}}(\mu_{2}(t))_{2,2}\frac{\operatorname{d}^{2}}{\operatorname{d}y^{2}}\mu_{1}(t)+\frac{\operatorname{d}^{2}}{\operatorname{d}x^{2}}\mu_{1}(t).$
@@
Now, for all $t\in[0,\log(M)]$, it holds $\rho({\operatorname{Cov}}(\mu_{1}(t)))=e^{t}\leq M$, and for all $\varphi\in\mathcal{C}_{c}^{\infty}(\mathbb{R}^{2})$,
$\int_{\mathcal{X}^{2,\infty}(\mathbb{R}^{2})}\mathcal{L}_{X}^{2}(\varphi)\operatorname{d}V[\mu_{1}(t)][X]$ (3.152)
$\qquad=\int_{\mathcal{X}^{2,\infty}(\mathbb{R}^{2})}X_{1}^{2}\operatorname{d}V[\mu_{1}(t)][X]\frac{\operatorname{d}^{2}}{\operatorname{d}x^{2}}\varphi+\int_{\mathcal{X}^{2,\infty}(\mathbb{R}^{2})}X_{2}^{2}\operatorname{d}V[\mu_{1}(t)][X]\frac{\operatorname{d}^{2}}{\operatorname{d}y^{2}}\varphi$
$\qquad+2\int_{\mathcal{X}^{2,\infty}(\mathbb{R}^{2})}X_{1}X_{2}\operatorname{d}V[\mu_{1}(t)][X]\frac{\operatorname{d}^{2}}{\operatorname{d}x\operatorname{d}y}\varphi$
$\qquad={\operatorname{Cov}}(\mu_{1}(t))_{1,1}\frac{\operatorname{d}^{2}}{\operatorname{d}x^{2}}\varphi+\Delta_{\rho({\operatorname{Cov}}(\mu_{1}(t)))}({\operatorname{Cov}}(\mu_{1}(t))_{2,2})\frac{\operatorname{d}^{2}}{\operatorname{d}y^{2}}\varphi.$

### v2

@@ 3.7.4 Strong Generalised Central Limit Theorem and uniqueness of smooth solutions
Let $\varepsilon>0$. First, Theorem 3.27 yields that there exists $W:[0,T]\mapsto\mathcal{W}_{p}(\mathcal{X}^{2,\infty}(M))$, which is $\frac{1}{p}$–Hölder–continuous with uniformly bounded (p+2)(p+2)–th moments and $\lambda>0$, such that for all $t\in[0,T]$, $\square^{W(t)}$ is $\lambda$–elliptic and $\mathcal{W}_{p}^{\infty}(V,W)\leq\varepsilon$. Since the $W(t)$ have uniformly bounded (p+2)(p+2)–th moments, one readily shows that for all $i,j\in\llbracket 0,n-1\rrbracket$,
@@
are uniformly in $W^{2,\infty}(M)$ for fixed $t$ and uniformly in $\mathcal{C}^{0,\frac{1}{p}}([0,T],\mathbb{R})$ for fixed $x$. Thus, (LMDE) has a unique distributional solution $s$ for $W$ with initial value $\mu_{0}$. Now $L:=\tfrac{\partial}{\partial t}-\square^{W(t)}$ is a second-order parabolic operator on $M\times(0,T]$, which is uniformly parabolic and the coefficients are uniformly $W^{2,\infty}$ in space and uniformly $\tfrac{1}{p}$–Hölder in time, so they are in $C^{2\alpha,\alpha}_{loc}(M\times(0,T])$ for $\alpha=\tfrac{1}{p}$. If $u\in\dot{\mathcal{B}}^{\prime}(M\times(0,T])$ ([21] where the decay is only spatially but uniformly in $[t,T]$ for all $t>0$) satisfies $Lu=0$ in the sense of distributions and $u(0)=\mu_{0}$, by the parabolic Schauder theory, one has
@@
Theorem 3.52 (Generalised Central Limit Theorem—Strong version).
Let $p>2$ and $T>0$. Let $\mu_{0}\in\mathcal{W}_{p}(M)$. Let $\bigl(\mathcal{}\mathfrak{P}(\frac{1}{N})\bigr)_{N\in\mathbb{N}}$ be a family of partitions of $[0,T]$. Assume:
@@
  • $V$ is a probability vector–field which satisfies:
    • There exists a constant $R>0$ such that for all $\mu\in\mathcal{W}_{p}(M)$, $V[\mu]\in\mathcal{W}_{p+2}(\mathcal{X}^{2,\infty}(M))$ and $\mathcal{M}_{p+2}(V[\mu])\leq R^{p+2}$.
@@
Then, (MDEd) has a unique smooth solution $\mu:[0,T]\to\mathcal{W}_{p}(M)$ (compare Definiton 3.30) with initial value $\mu_{0}$ and the Average Flow Approximation Series $\bigl(\mu^{\mathcal{}\mathfrak{P}(\frac{1}{N})}\bigr)_{N\in\mathbb{N}}$ (see Definition 3.29) converges uniformly to $t\mapsto\mu(t)$ in $\mathcal{W}_{p}(M)$. Furthermore, $t\mapsto\mu(t)$ is $\frac{1}{p}$–Hölder–continuous.
@@
Let $\mu_{0}\in\mathcal{C}^{\infty}_{c}(M)$. Proposition 3.51 yields that for all $\frac{1}{p}$–Hölder–continuous $\mu:[0,T]\to\mathcal{W}_{p}(M)$ with $\mu(0)=\mu_{0}$, the linear MDE (LMDE) has a unique smooth solution for $t\mapsto V[\mu(t)]$ with initial value $\mu_{0}$ in $\mathcal{W}_{p}(M)$. Hence, Theorem 3.48 yields that (MDEd) has a unique smooth solution with initial value $\mu_{0}$ in $\mathcal{W}_{p}(M)$. Finally, we extend the uniqueness on arbitrary $\mu_{0}\in\mathcal{W}_{p}(M)$ with the last point of Theorem 3.46. Now, the rest of the statement is a consequence of Theorem 3.39. $\square$
@@
Remark 3.54.
Theorem 3.52 still holds if we slightly relax the spatial regularity of our vector–field probabilities. For all the above proofs, it is sufficient to require our tangent vectors $V$ to be supported in $\mathcal{X}^{1,\infty}(M)$, such that for all $X\in{\operatorname{supp}}(V)$, $X-\overline{V}\in\mathcal{X}^{2,\infty}(M)$. In other words, the centered vector–field probability $\tau^{-\overline{V}}_{*}V\in\mathcal{W}_{p}(\mathcal{X}^{2,\infty}(M))$, while we only assume $W^{1,\infty}$–regularity for the average vector–field $\overline{V}$. In fact, while some of the above results fail in all generality if we weaken these assumptions even further, Theorem 3.52 still holds if we only assume $\tau^{-\overline{V}}_{*}V\in\mathcal{W}_{p}(\mathcal{X}^{1+\delta,\infty}(M))$ for some $\delta>0$.
Remark 3.55.
Smooth solutions of (MDEd) are particularly relevant for applications, since they are predictable. Knowing $\mu_{0}$ and $V[\mu_{0}]$ is sufficient to compute an approximation of $\mu(t)$, for small $t>0$, by the means of AFAS. To approximate any non–smooth solution of (MDEd), even for $t>0$ arbitrarily small, one needs to know $V$ on a whole neighborhood of $\mu_{0}$ in $\mathcal{W}_{p}(M)$.
@@ 3.7.5 A counter–example to general uniqueness of solutions
General solutions of (MDEd) are however not unique in general, even if $V$ is Lipschitz and the $\square^{V}_{\mu}$ are uniformly elliptic for all $\mu\in\mathcal{W}_{p}(M)$. Let us give an example of two distinct solutions to (MDEd) with same initial value.
@@
There exists a probability vector–field $V$ which is Lipschitz with respect to the pp–Wasserstein distance and such that $\square^{V}_{\mu}$ is strongly elliptic for all $\mu\in\mathcal{W}_{p}(M)$, such that solutions to (MDEd) are not unique.
@@
For any $r\in\mathbb{R}$ and $v_{1},v_{2}\in\mathbb{R}^{2}$ with $\bigl<v_{1},v_{2}\bigr>=0$, denote by $\mathcal{C}_{r}$ the circle with radius $r$ and by $\mathcal{E}_{v_{1},v_{2}}$ the ellipse with semimajor axis $v_{1}$ and semiminor axis $v_{2}$. Finally, define the probability vector–field
$V:\mathcal{W}_{p}(\mathbb{R}^{2})$ $\mapsto\mathcal{W}_{p}(\mathcal{X}^{2,\infty}(\mathbb{R}^{2}))$
@@
where, by abuse of notation, for all $X\in\mathbb{R}^{2}$, we identify the constant vector–field $v_{X}:y\mapsto X$ with $X$ here. Now, we show that $V$ is locally Lipschitz with respect to the pp–Wasserstein distance around $\mu_{0}$, where $\mu_{0}$ is the standard normal distribution. Using Proposition 2.15, there exist $R>0$, $B>0$ and $b>0$, such that for all $\mu\in B(\mu_{0},R)$, $r({\operatorname{Cov}}(\mu))\geq b$ and $\rho({\operatorname{Cov}}(\mu))\leq B$. Thus, there exist $C>0$, such that for all $\rho,r_{1},r_{2}\in[b,B]$, $\bigl|\Delta_{\rho}(r_{1})-\Delta_{\rho}(r_{2})\bigr|\leq C|r_{1}-r_{2}|$. Now, Proposition 2.15 yields that $V$ is locally Lipschitz around $\mu_{0}$. Furthermore, $V$ is uniformly bounded by $M>0$.
Now, we show that there exist two distinct solutions to (MDEd) with initial value $\mu_{0}$. Let $T>0$. Define
$\mu_{1}:[0,T]\mapsto\biggl((x,y)\mapsto\frac{1}{2\pi\sqrt{(2t+1)e^{2t}}}e^{-\frac{1}{2}\bigl(\frac{x^{2}}{2t+1}+\frac{y^{2}}{e^{2t}}\bigr)}\biggr)$ (3.157)
$\mu_{2}:[0,T]\mapsto\biggl((x,y)\mapsto\frac{1}{2\pi\sqrt{(2t+1)e^{2t}}}e^{-\frac{1}{2}\bigl(\frac{y^{2}}{2t+1}+\frac{x^{2}}{e^{2t}}\bigr)}\biggr).$
@@
$\frac{\operatorname{d}}{\operatorname{d}t}\mu_{1}(t)=\frac{\operatorname{d}^{2}}{\operatorname{d}x^{2}}\mu_{1}(t)+{\operatorname{Cov}}(\mu_{1}(t))_{2,2}\frac{\operatorname{d}^{2}}{\operatorname{d}y^{2}}\mu_{1}(t),$ (3.158)
$\frac{\operatorname{d}}{\operatorname{d}t}\mu_{2}(t)=\frac{\operatorname{d}^{2}}{\operatorname{d}y^{2}}\mu_{2}(t)+{\operatorname{Cov}}(\mu_{2}(t))_{1,1}\frac{\operatorname{d}^{2}}{\operatorname{d}x^{2}}\mu_{2}(t).$
@@
Now, for all $t\in[0,\sqrt{\log(M)}]$, it holds that $\rho({\operatorname{Cov}}(\mu_{2}(t)))=e^{2t}\leq M$, and for all $\varphi\in\mathcal{C}_{c}^{\infty}(\mathbb{R}^{2})$,
$\int_{\mathcal{X}^{2,\infty}(\mathbb{R}^{2})}\mathcal{L}_{X}^{2}(\varphi)\operatorname{d}V[\mu_{2}(t)](X)$ (3.159)
$\hskip 18.49988pt=\int_{\mathcal{X}^{2,\infty}(\mathbb{R}^{2})}X_{1}^{2}\operatorname{d}V[\mu_{2}(t)](X)\frac{\operatorname{d}^{2}}{\operatorname{d}x^{2}}\varphi+\int_{\mathcal{X}^{2,\infty}(\mathbb{R}^{2})}X_{2}^{2}\operatorname{d}V[\mu_{2}(t)](X)\frac{\operatorname{d}^{2}}{\operatorname{d}y^{2}}\varphi$
$\hskip 18.49988pt+2\int_{\mathcal{X}^{2,\infty}(\mathbb{R}^{2})}X_{1}X_{2}\operatorname{d}V[\mu_{2}(t)](X)\frac{\operatorname{d}^{2}}{\operatorname{d}x\operatorname{d}y}\varphi$
$\hskip 18.49988pt={\operatorname{Cov}}(\mu_{2}(t))_{1,1}\frac{\operatorname{d}^{2}}{\operatorname{d}x^{2}}\varphi+\Delta_{\rho({\operatorname{Cov}}(\mu_{2}(t)))}({\operatorname{Cov}}(\mu_{2}(t))_{2,2})\frac{\operatorname{d}^{2}}{\operatorname{d}y^{2}}\varphi.$